Circular No. 6686

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SCORPII 1997 R. W. Argyle, Royal Greenwich Observatory, reports an accurate opti-

cal position for N Sco 1997 obtained from 9 observations with the Carlsberg Automatic Meridian Circle on La Palma: α = 17^{h}54^{m}11^{s}.203 ± 0^{s}.006, δ = −30^{o}02′53′′.17 ± 0′′.11 (equinox J2000.0, Hipparcos reference frame, epoch 1997.46). The following V magnitudes (±0.08 mag) were obtained: June 15.06 UT, 11.95; 16.06, 12.14; 17.06, 12.18; 18.06, 12.26; 19.05, 12.45; 22.04, 12.76; 23.04, 12.43.

COMET C/1995 O1 (HALE-BOPP) M. Sakamoto, Misato Observatory; B. Suzuki, Misato Technological

High School; H. Kawakita, Osaka; R. Furusho, Kobe University; N. Yamamoto, Science University of Tokyo; and J. Watanabe, National Astronomical Observatory, communicate: “Neutral-sodium emission images of comet C/1995 O1 were obtained since Feb. 20, using a narrow-band filter set that was designed for neutral-sodium emission at 589 nm at the Misato 1.05-m reflector (+ liquid-N_{2}-cooled CCD camera). The size of the images correspond to 500 000 km × 500 000 km. The most remarkable phenomenon is the ‘sodium jet’ structure released in the solar direction from the nucleus. The sodium distribution of the near-nucleus region resembles the dust profile, though the sodium jet slips off of the dust jet in the antisolar direction. In a similar way, the Na distribution is more sensitive than is the dust tail to radiation pressure. More detailed information can be

found at http://www.obs.misato.wakayama.jp/HB/index-e.html.”

COMET C/1997 J2 (MEUNIER-DUPOUY) K. Churyumov, Kiev University Observatory, reports that he and V.

Vlassyuk obtained two spectrograms (range 460–550 nm) of comet C/1997 J2 on June 20.95 UT with the Special Astrophysical Observatory’s 6-m telescope (+ Planetary spectrograph with long slit and CCD) at Mount Pastukhov; the spectra show solar continuum and noticeable C_{3} emission.

COMET 2P/ENCKE Visual m_{1} estimates: June 1.34 UT, 6.2 (D. A. J. Seargent, Cowra,

N.S.W., Australia, 25×100 binoculars); 9.43, 6.5 (A. Pearce, City Beach, W. Australia, 0.2-m reflector); 23.45, 7.3 (Pearce, 20×80 binoculars).

1997 June 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green